CAN SOMEONE PLEASE EXPLAN FO ME WHATS WRONG WITH MY CODE WHEN EVEN I RUN IT IT GIVE ME THIS MASSAGE IM USING python BTW

Traceback (most recent call last):                                                                                                                                              

  File "../resource/asnlib/public/RUN.py", line 20, in <module>                                                                                                                 

    exec(source_code, dict())                                                                                                                                                   

  File "<string>", line 60, in <module>                                                                                                                                         

NameError: name 'total_rough_sod' is not defined )

from math import pi

length = float(input('Enter Course Length:'))
width = float(input('Enter Course Width:'))


def calculate_smooth_sod(width):
grean_area_raduis= (width/2)/2
total_smooth_sod=2*pi*(grean_area_raduis**2)
return total_smooth_sod

def calculate_sand_trap_area(width):
sand_trap_radius=(width/3)/2
sand_trap_area=pi*(sand_trap_radius**2)
return sand_trap_area

def calculate_rough_sod(length, width, total_smooth_sod, sand_trap_area):
total_area= length*width
total_rough_sod= total_area-total_smooth_sod-sand_trap_area
return total_rough_sod

def calculate_bricks(width):
length_of_brick_wall= pi*(width*3/6)
number_of_bricks= length_of_brick_wall*3
return number_of_bricks

def calculate_sand(width):
width_in_ft= width*3
sand_trap_radius= width_in_ft/6
sand_trap_area=pi*(sand_trap_radius**2)
total_sand=(sand_trap_area*100)/2000
return total_sand

def calculate_bushes(length, width):
perimeter=2*(length+width)
total_bushes=perimeter-2
return total_bushes

def calculate_total_time(total_rough_sod, total_smooth_sod):
total_time=(total_rough_sod*0.5)+(total_smooth_sod)
return total_time

60-print("Total square yards of rough sod: " + str(total_rough_sod))
61-print("Total square yards of smooth sod: " + str(total_smooth_sod))
62-print("Tons of sand: " + str(total_sand))
63-print("Number of retaining wall bricks: " + str(number_of_bricks))
64-print("Number of bushes: " + str(total_bushes))
65-print("Total mowing time (mins): " + str(total_time))
Answer #1
#variables declared inside functions are not accessible outside. Also, you never called the created functions on the values inputted
from math import pi

def calculate_smooth_sod(width):
    grean_area_raduis= (width/2)/2
    total_smooth_sod=2*pi*(grean_area_raduis**2)
    return total_smooth_sod

def calculate_sand_trap_area(width):
    sand_trap_radius=(width/3)/2
    sand_trap_area=pi*(sand_trap_radius**2)
    return sand_trap_area

def calculate_rough_sod(length, width, total_smooth_sod, sand_trap_area):
    total_area= length*width
    total_rough_sod= total_area-total_smooth_sod-sand_trap_area
    return total_rough_sod

def calculate_bricks(width):
    length_of_brick_wall= pi*(width*3/6)
    number_of_bricks= length_of_brick_wall*3
    return number_of_bricks

def calculate_sand(width):
    width_in_ft= width*3
    sand_trap_radius= width_in_ft/6
    sand_trap_area=pi*(sand_trap_radius**2)
    total_sand=(sand_trap_area*100)/2000
    return total_sand

def calculate_bushes(length, width):
    perimeter=2*(length+width)
    total_bushes=perimeter-2
    return total_bushes

def calculate_total_time(total_rough_sod, total_smooth_sod):
    total_time=(total_rough_sod*0.5)+(total_smooth_sod)
    return total_time

length = float(input('Enter Course Length: '))
width = float(input('Enter Course Width: '))

print("Total square yards of rough sod: " + str (calculate_rough_sod(length,width,calculate_smooth_sod(width),calculate_sand_trap_area(width))))
print("Total square yards of smooth sod: " + str(calculate_smooth_sod(width)))
print("Tons of sand: " + str(calculate_sand(width)))
print("Number of retaining wall bricks: " + str(calculate_bricks(width)))
print("Number of bushes: " + str(calculate_bushes(length,width)))
print("Total mowing time (mins): " + str(calculate_total_time(calculate_rough_sod(length,width,calculate_smooth_sod(width),calculate_sand_trap_area(width)),calculate_smooth_sod(width))))